INAF queue observing, 2016 Jan 27-28

Observer: F. Cusano, A. Rossi
Telescope Operator: S. Allanson
Support Astronomer: Olga Kuhn
Instrument: LUCI/MODS

Summary

At the sunset the sky is clear.

Night duration (12\xB0 el twilight) 19:20-6:10 ( 02:20-13:10 UT) = 10 h 50m

Observing time: * 10h* 30m (97%)

Weather time loss: 0 h (0%)

technical time loss * 0h* 20m (3%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
MODScalib        
Palazzi Eliana1 LUCI zJ,HKspec+LBC UBVr 2.7h  
Gilli id3605 zJ 2.5h  
DeBarros M103GLASS2c zJ 2h  

Details

00:05 MODS slitless flat red grating mods1r.20160128.0002-000

01:41 starting Palazzi program
  • acquisition telluric luci2.20160128.0001-04
  • seeing 1.0" on the guider at 1:43
  • spec telluric luci2.20160128.0005-13

02:05 slew on Palazzi target Eliana1: parallel observations LUCI2 spectra (zJspec,HKspec) + LBC-blue sequence of UBVr_sloan
  • acquisition luci2.20160128.0014-17
  • seeing 0.8" on the guider at 02:11
  • spectra luci2.20160128.0018-23 zJspec
  • seeing 0.8" on the guider at 02:34
  • simultaneous obs with LBC-blue: waiting for dofpia to finish at 02:36
  • 02:39 start LBC-blue Eliana1U: not good due to luci offset
  • 02:40 problem with scheduler.
  • 02:46 start LUCI ob again, LBC-blue Eliana1B, 1V, 1r
  • keep sequence Eliana1 UBVr obs with LBCs
  • seeing 0.7 at 02:50 on LBC-b
  • seeing 0.6 at 3:05 on guider
  • 3:17 spectra luci2.20160128.0024-0029 HKspec
  • seeing 0.6 at 3:20 on guider
  • 3:53 spectra luci2.20160128.0030-0035 zJ
  • 4:29 spectra luci2.20160128.0036-41 HK
  • problem with scheduler 10 min lost 0037 is outside slit. we apllied manual offset +60" then should be fine
  • problem LBC mirror. no LBC images between 4:35 -5:20 image 05408 052822 not good
  • 5:14 spectra luci2.20160128.0042-47 zJ
  • dofpia worked LBC with PA=270 to help mirror movement. First image useful is 053946 but is rotated (270deg).
  • rotation does not help, we use PA=0 LBC obs
  • seeing 0.9 at 5:45 on guider
  • 6:00 spectra luci2.20160128.0048-49 HK are ok then scheduler problem
  • 6:10 stop observations

06:15 start telluric for Palazzi again (different airmass)
  • acquisition telluric luci2.20160128.0050-52
  • seeing 0.7" on the guider at 06:20
  • spec telluric luci2.20160128.0053-56
  • spec telluric luci2.20160128.0057-60

06:35 start telluric for Gilli
  • acquisition telluric luci2.20160128.0061-63
  • check pointing
  • spec telluric luci2.20160128.0064-0067

06:50 slew to science of Gilli
  • acq luci2.20160128.0068-0071
  • spec luci2.20160128.0072-0101
  • seeing 0.7" on the guider at 7:00
  • we start dofpia with LBC-B at 7:01
  • starting imaging LBC-B, we run 1_LBC_3605_r_short.ob
  • seeing 0.6" on the guider at 7:32
  • dofpia 8:01
  • seeing 0.4" on the guider at 8:25
  • dofpia 8:45
  • seeing 0.6 " at 9:00

09:56 slew to telluric of Gilli
  • preset failed, the TO checks the pointing
  • acq luci2.20160128.0102-105
  • spec luci2.20160128.0106-0108
  • seeing 0.5" at 10:00

10:18 slew to telluric of de Barros
  • acq luci2.20160128.0109-0111
  • spec luci2.20160128.0112-0119
  • seeing 0.5" on the guider

10:40 slew to science of de Barros
  • acq luci2.20160128.0120-0123
  • seeing 0.4" at 10:43
  • spectra luci2.20160128.0124-0147
  • seeing 0.6" at 11:17
  • seeing 0.4" at 11:41
  • finish at 13:17

 
Topic revision: r20 - 01 Feb 2016, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback